# Data-Efficient Fine-Tuning of Vision-Language Models for Diagnosis of Alzheimer’s Disease

### v1

@@ 4.3.1 Impact of individual component
We assess the effectiveness of each component in our fine-tuning pipeline on the ADNI dataset using M3D as the backbone. As shown in Table 3, compared to the baseline (a), fine-tuning the image encoder (b) or the text encoder (c) improves classification accuracy, with the highest accuracy achieved when fine-tuning both encoders (d). Moreover, we find that reports generated from metadata provide the largest performance gain during fine-tuning, highlighting the significant value and potential of incorporating metadata.
4.3.2 Impact of fine-tuning sample size
500 1000 1500 2000
ACC 63.1% 66.8% 78.8% 81.3%
Table 5: Classification performance with varying fine-tuning sample size on ADNI dataset
4.3.3 Impact of fine-tuning strategies
@@ 4.3.4 Interpreting biomarker importance
Stage Top 1 Top 2 Top 3
NC vol entorhinal size
MCI vol ventricular entorhinal
AD brain entorhinal hippocampal
Table 6: Top 3 contributing tokens at each disease progression stage based on attribution scores from our model interpretation for the same subject. Top 1, 2, and 3 indicate the three most important contributing biomarkers.
To better understand how our model utilizes textual features for image-text alignment, we adopt an attribution method based on Integrated Gradients [36] to interpret the contribution of individual text tokens to our model’s matching scores. As shown in Table 6, for a single subject across three disease progression stages, the top contributing words shift notably. At the earliest stage, “vol”,“entorhinal”, and “size" show high attribution, indicating the model’s focus on volumetric measures. In the intermediate stage, the attribution to “entorhinal”, “ventricular”, and related biomarkers receive high attribution, reflecting advancing structural changes. By the latest stage, tokens such as “brain" and “hippocampal" become more prominent among the top contributors, indicating a shift in the model’s focus potentially related to atrophy or disease progression. This dynamic pattern in attribution scores highlights that specific biomarkers such as the “entorhinal" and “ventricular" are critical indicators across disease progression. These results demonstrate that our model’s decisions are closely related to key clinical biomarkers, supporting its relevance and reliability in disease assessment.
@@ 4.3.5 Visualization analysis
When using whole brain volume as the only input, the attention tends to spread across broader brain areas. This suggests that whole brain volume may introduce less specific guidance for localization, and the model’s attention becomes more diffuse. Overall, these results highlight that certain biomarkers—particularly entorhinal volume and ventricular size—are more informative and effective in guiding the model toward disease-relevant visual features.
@@ Acknowledgment
Data for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904). OASIS-2 data was obtained from the Open Access Series of Imaging Studies (OASIS) project.

### v3

@@ 4.3.4 Interpreting biomarker importance
To understand how our model leverages textual features for image-text alignment, we use Integrated Gradients [37] to attribute contributions of individual text tokens to matching scores. As shown in Figure 2, for a single subject across three disease stages, the top contributing words shift: early-stage tokens like “vol”, “entorhinal”, and “size” highlight volumetric measures; intermediate-stage tokens such as “entorhinal” and “ventricular” reflect advancing structural changes; and late-stage tokens like “brain” and “hippocampal” indicate focus on atrophy or progression. These results show that the model’s decisions align with key clinical biomarkers, supporting its relevance in disease assessment.
Figure 2: Heatmap of integrated gradients showing the importance of each biomarker for different disease progression stages (darker color indicates higher importance).
@@ 4.3.5 Visualization analysis
When using whole brain volume as the only input, the attention tends to spread across broader brain areas. This suggests that whole brain volume may introduce less specific guidance for localization, and the model’s attention becomes more diffuse. Overall, these results highlight that certain biomarkers – particularly entorhinal volume and ventricular size – are more informative and effective in guiding the model toward disease-relevant visual features.
@@ Acknowledgment
Data for this project was funded by the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (National Institutes of Health Grant U01 AG024904). OASIS-2 data was obtained from the Open Access Series of Imaging Studies (OASIS) project, and AIBL data was obtained from the Australian Imaging, Biomarkers and Lifestyle (AIBL) study.